OpenStudy (anonymous):

A mule deer can run one over four of a mile in 25 seconds. At this rate, which expression can be used to determine how fast a mule deer runs in miles per hour? twenty-five seconds over one-fourth mile = 36 miles per hour one-fourth mile over twenty-five seconds = 36 miles per hour one-fourth mile over twenty-five seconds = 100 miles per hour twenty-five seconds over one-fourth mile = 100 miles per hour

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

hmm, what do you think texaschick101?

OpenStudy (anonymous):

thanks haydenhd

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@texaschic101 what do you think?

OpenStudy (anonymous):

do you agree with haydenhd?

OpenStudy (anonymous):

help texaschic101

OpenStudy (anonymous):

@phi

OpenStudy (texaschic101):

I am not sure...sorry

OpenStudy (anonymous):

oh ok. its allright though. thanks for helping!!!!!!!!!!!! :D

OpenStudy (anonymous):

i might be wrong cause i dont think any animal could run 100 mph

OpenStudy (anonymous):

hahaha WOW :D

OpenStudy (anonymous):

theyre probably making it up then

OpenStudy (anonymous):

maybe...

OpenStudy (anonymous):

its either 36 mph or 100 mph

OpenStudy (anonymous):

maybe A?

OpenStudy (anonymous):

@angel12310

OpenStudy (texaschic101):

1/4 of a mile in 25 seconds = 1 mile in 100 seconds and there is 3600 seconds per hr 1/100 = x/3600 100x = 3600 x = 36 miles per hr ??? I am thinking A or B....not sure which one

OpenStudy (anonymous):

ok. so it all depends on the fraction

OpenStudy (anonymous):

prob a

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

nope, not a

OpenStudy (anonymous):

Can you reset the test

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Then put B

OpenStudy (anonymous):

ok lemme try

OpenStudy (anonymous):

It would be stupid if a deer could run 100 miles an hour!

OpenStudy (texaschic101):

that deer is hauling retrice..lol

OpenStudy (anonymous):

well look, i have another q. lemme open another browser thingy

OpenStudy (texaschic101):

oops...they replaced my word...lol

OpenStudy (anonymous):

Was the answer B right?

OpenStudy (anonymous):

look, i have other questions to finish in the same test so i still havent tried

OpenStudy (anonymous):

phi what do you think is the answer?

OpenStudy (phi):

you want to do miles/hours you are given 1/4 miles you are given 25 seconds seconds is time, but you have to change 25 seconds into hours the way to do that is 25 seconds * 1 hour/3600 seconds = 25/3600 hours simplify. flip multiply times 1/4

OpenStudy (anonymous):

25/3600 times 1/4?

OpenStudy (phi):

remember we are doing 1/4 *divided* by 25/3600 flip the "bottom fraction" and then multiply

OpenStudy (anonymous):

that makes sense which means B s the answer.

OpenStudy (anonymous):

oh. so then the answer is b. lemme try

OpenStudy (anonymous):

Which unit rate is equivalent to 17 miles per gallon? three gallons over fifty-one miles fifty-one miles over three gallons thirty-four miles over four gallons four gallons over thirty-four miles

OpenStudy (anonymous):

i have to answer that question to submit the test

OpenStudy (phi):

First, which choices have the correct order? do you miles / gallons or gallons/ miles ?

OpenStudy (anonymous):

gallons/miles

OpenStudy (anonymous):

so then its either a or d

OpenStudy (phi):

equivalent to 17 miles per gallon? do you see MILES PER GALLON and the PER means "divided by"

OpenStudy (texaschic101):

this can be solved with proportions

OpenStudy (anonymous):

oops miles over gallons

OpenStudy (anonymous):

b or c

OpenStudy (anonymous):

texaschic101?

OpenStudy (phi):

now check each choice to see which gives 17

OpenStudy (texaschic101):

17/1 = 51/3 (cross multiply and if it comes out equal, this is your answer) (51)(1) = (17)(3) 51 = 51 (correct) 17 miles per gallon = 51 miles per 3 gallons

OpenStudy (anonymous):

ok thanks :D

OpenStudy (texaschic101):

Phi probably had an easier way to solve this

OpenStudy (anonymous):

A